# Ground-state phase diagram of $S=1/2$ Heisenberg model on 2D square-hexagon-octagon lattice

### v1

@@ IV CONCLUSIONS
In this paper, we have investigated the ground-state properties and critical behavior of the $S=1/2$ Heisenberg model on the SHO lattice. The SHO lattice, characterized by a polygonal nested structure with periodically arranged four-, six-, and eight-membered rings on a two-dimensional plane, offers a rich landscape for exploring quantum magnetic phenomena. To study the rich ground-state phases of the Heisenberg model on the SHO lattice, we introduced three types of interactions: two nearest-neighbor interactions $J_{1}$ and $J_{2}$ and a selected third nearest-neighbor interaction $J_{3}$ along $x$ direction. Our numerical approach combined the strengths of SSE-QMC and DMRG methods, with results analyzed through finite-size scaling and extrapolation techniques. This comprehensive analysis enabled us to construct a detailed phase diagram of the ground state, identifying five distinct phases: the AFM phase, hexagon phase, OSD phase, LSD phase and Haldane-like SPT phase. We focused on representative paths within the phase diagram to present and analyze key data. Especially for the Haldane-like symmetry-protected topological (SPT) phase, we employed the DMRG method to identify its edge states and topological properties. By examining the degeneracy of the ground-state energy levels and analyzing the entanglement spectrum under open boundary conditions, we were able to characterize the unique nature of this phase. Moreover, by employing the finite-size scaling analysis method, we obtained the critical exponent $\nu$ for the phase transitions in the SHO lattice. Comparison with the known critical exponent of the 3D $O(3)$ universality class revealed that all phase transitions from non-magnetic phases to AFM phase in the SHO lattice Heisenberg model belong to this universality class.
Our work not only advances the theoretical understanding of quantum spin systems on complex lattices but also holds practical implications. Once the lattice magnetic materials with a SHO structure (which may not have strong $J_{3}$ interactions) and Heisenberg interactions be successfully synthesized in the future, our numerical results will provide a benchmark for comparing and verifying experimental measurements in real physical systems. From a fundamental theoretical standpoint, our work not only provides a platform for exploring complex phase behavior but also delivers robust, precise numerical results.
@@ Acknowledgements.
We would like to thank Zheng-Xin Liu for the helpful discussions. This project is supported by NKRDPC-2022YFA1402802, NSFC-12474248, NSFC-92165204, NSFC-12494591, Leading Talent Program of Guangdong Special Projects (Grant No. 201626003), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515120013), Guangdong Fundamental Research Center for Magnetoelectric Physics (Grant No.2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (Grant No.GDZX2401010).

### v2

@@ IV CONCLUSIONS
In this paper, we have investigated the ground-state properties and critical behavior of the $S=1/2$ Heisenberg model on the SHO lattice. The SHO lattice, characterized by a polygonal nested structure with periodically arranged four-, six-, and eight-membered rings on a two-dimensional plane, offers a rich landscape for exploring quantum magnetic phenomena. To study the rich ground-state phases of the Heisenberg model on the SHO lattice, we introduced three types of interactions: two nearest-neighbor interactions $J_{1}$ and $J_{2}$ and a selected third-nearest-neighbor interaction $J_{3}$ along the $x$ direction. Our numerical approach combined the strengths of SSE-QMC and DMRG methods, with results analyzed through finite-size scaling and extrapolation techniques. This comprehensive analysis enabled us to construct a detailed phase diagram of the ground state, identifying five distinct phases: the AFM phase, hexagon phase, OSD phase, LSD phase, and Haldane-like SPT phase. We focused on representative paths within the phase diagram to present and analyze key data. Especially for the Haldane-like SPT phase, we also employed the DMRG method to identify its edge states and topological properties. By examining the degeneracy of the ground-state energy levels and analyzing the entanglement spectrum under open boundary conditions, we were able to characterize the unique nature of this phase. To connect our findings to real materials, we also investigated the robustness of the SPT phase against other possible interactions. Moreover, by employing the finite-size scaling analysis method, we obtained the critical exponent $\nu$ for the phase transitions in the SHO lattice. Comparison with the known critical exponent of the 3D $O(3)$ universality class revealed that all phase transitions from nonmagnetic phases to the AFM phase in the SHO lattice Heisenberg model belong to this universality class.
Our work, with the Haldane-like SPT phase as one of the important results, significantly advances the theoretical understanding of topological states in quantum spin systems. This finding is of fundamental importance as it provides a platform for exploring complex phase behavior and delivers robust, precise numerical results on a nontrivial lattice.
@@ Acknowledgements.
We would like to thank Zheng-Xin Liu for the helpful discussions. This project is supported by NSFC-12474248, NKRDPC-2022YFA1402802, NSFC-92165204, NSFC-12494591, Leading Talent Program of Guangdong Special Projects (Grant No. 201626003), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515120013), Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (No. 2022B1212010008), Guangdong Fundamental Research Center for Magnetoelectric Physics (Grant No. 2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2401010).